OpenStudy (anonymous):

how do i evaluate this expression: x(x + y) when x=-2 1/2 and y= 3 3/4?

OpenStudy (bahrom7893):

just plug them in.

OpenStudy (bahrom7893):

(-5/2)*((-5/2)+(15/4))

OpenStudy (bahrom7893):

come on, you can finish the rest...

OpenStudy (bahrom7893):

(-5/2)*((15/4) - (10/4))

OpenStudy (bahrom7893):

(-5/2)*(5/4) = -25/8

OpenStudy (anonymous):

mikevucina do you get how to do it?

OpenStudy (anonymous):

i do know

OpenStudy (anonymous):

now

OpenStudy (anonymous):

do you know the order of operations?

OpenStudy (anonymous):

ye

OpenStudy (anonymous):

what comes first parentheses or multiplication? pop quiz lol

OpenStudy (anonymous):

parentheses

OpenStudy (anonymous):

nice

OpenStudy (anonymous):

okay 3x squared when x= -1.5 is 6.25?

OpenStudy (anonymous):

what comes first, exponents or multiplication?

OpenStudy (anonymous):

exponents

OpenStudy (anonymous):

btw im just learning this myself with you

OpenStudy (anonymous):

true

OpenStudy (anonymous):

bahrom do you square the integer first and then multiply by 3?

OpenStudy (anonymous):

square number first then integer i think

OpenStudy (anonymous):

i think you square the x (which is -1.5 and then multiply by 3

OpenStudy (anonymous):

unless its (3x)squared

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

because parentheses comes before exponents

OpenStudy (bahrom7893):

sorry where?

OpenStudy (anonymous):

3x squared when x= -1.5

OpenStudy (bahrom7893):

depends on the parenthesis

OpenStudy (anonymous):

question is 3x squared + 5x - 12 when x= -1.5

OpenStudy (anonymous):

are there parentheses?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

so exponents come first, -1.5^2 = what?

OpenStudy (anonymous):

2.25

OpenStudy (anonymous):

so now rewrite the equation

OpenStudy (anonymous):

with the new piece of info

OpenStudy (anonymous):

6.75 + 5x - 12

OpenStudy (anonymous):

boom, and we can go ahead and do the 5x, its 5 * 2.25

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

i mean 5 *x or 5* -1.5

OpenStudy (anonymous):

can you do that and then rewrite the equation with the new info?

OpenStudy (anonymous):

the answer should be 19.25 but i got 11.25?

OpenStudy (anonymous):

hmmm, guess we need help, im asking bahrom

OpenStudy (bahrom7893):

can you guys draw out the actual question?

OpenStudy (anonymous):

question is 3x2 + 5x - 12 where x = -1.5

OpenStudy (anonymous):

3x^2

OpenStudy (bahrom7893):

3*(-1.5)^2+5*(-1.5)-12= 3*(2.25)-5*(1.5)-12

OpenStudy (anonymous):

howd you make it -5 in the second line?

OpenStudy (bahrom7893):

6.75-7.5-12= -12.5

OpenStudy (bahrom7893):

5*(-1.5) that negative in front of -1.5 makes the whole thing negative

OpenStudy (anonymous):

answer is supposed to be 19.25

OpenStudy (anonymous):

ok, so that looks good. mike youre saying that the answer should be 19.25? thats weird, bahrom looks right to me, i mean it makes sense

OpenStudy (bahrom7893):